Tier1 Operations Report 2018-04-04

From GridPP Wiki
Jump to: navigation, search

RAL Tier1 Operations Report for 4th April 2018

Review of Issues during the week 27th March to 4th April 2018
  • A relatively quiet week over the Easter weekend. There was a problem with CMS (Castor) on 28th March.
Current operational status and issues
  • The problems with data flows through the site firewall being restricted is still present. New firewall equipment has been delivered and is scheduled for installation in a few weeks time.
Resolved Castor Disk Server Issues
  • None.
Ongoing Castor Disk Server Issues
  • None
Limits on concurrent batch system jobs.
  • CMS Multicore 550
Notable Changes made since the last meeting.
  • Four disk servers have been deployed into Castor GenTape and the older ones withdrawn from service.
  • New Docker images rolled out across entire batch farm. (These fix a security issue with Singularity).
  • Rolled out CMS xrootd re-direction on Echo external gateways - to match that in the xrootd re-directors on the Worker Nodes.
  • At the time of the meeting an update is being applied to Echo which patches the 'backfill' bug.
Entries in GOC DB starting since the last report.
  • None
Declared in the GOC DB
  • None.
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.
  • Replacement (upgrade) of RAL firewall scheduled for a few weeks time.

Listing by category:

  • Echo
    • Applying minor CEPH update to fix the "backfill" bug. Increase number of placement groups and add more capacity.
  • Castor:
    • Update systems to use SL7 and configured by Quattor/Aquilon. (Tape servers done)
    • Move to generic Castor headnodes.
  • Networking
    • Extend the number of services on the production network with IPv6 dual stack. (Done for Perfsonar, FTS3, all squids and the CVMFS Stratum-1 servers).
    • Replacement (upgrade) of RAL firewall.
  • Internal
    • DNS servers will be rolled out within the Tier1 network.
  • Infrastructure
    • Testing of power distribution boards in the R89 machine room is being scheduled for some time late July / Early August. The effect of this on our services is being discussed.
Open GGUS Tickets (Snapshot during morning of meeting)
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
117683 none on hold less urgent 18/11/2015 28/03/2018 Information System CASTOR at RAL not publishing GLUE 2 EGI
124876 ops on hold less urgent 07/11/2016 13/11/2017 Operations [Rod Dashboard] Issue detected : hr.srce.GridFTP-Transfer-ops@gridftp.echo.stfc.ac.uk EGI
127597 cms on hold urgent 07/04/2017 20/03/2018 File Transfer Check networking and xrootd RAL-CERN performance EGI
132589 lhcb waiting for reply very urgent 21/12/2017 03/04/2018 Local Batch System Killed pilots at RAL WLCG
133992 atlas in progress less urgent 12/03/2018 15/03/2018 File Transfer RAL-LCG2-ECHO: No such file or directory EGI
134136 atlas in progress less urgent 20/03/2018 27/03/2018 File Transfer RAL-LCG2: [SE][Ls][SRM_INVALID_PATH] No such file or directory EGI
134342 cms in progress urgent 29/03/2018 03/04/2018 CMS_Data Transfers CERN->MSS rate test WLCG
GGUS Tickets Closed Last week
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
133719 atlas closed urgent 27/02/2018 28/03/2018 File Transfer Transfers to RAL-LCG2-ECHO failing WLCG
133752 atlas closed very urgent 01/03/2018 28/03/2018 File Transfer RAL FTS service appears broken WLCG
133764 snoplus.snolab.ca closed very urgent 01/03/2018 28/03/2018 Information System BDII missing SFU information EGI
134030 other solved less urgent 14/03/2018 03/04/2018 Other request to replicate larsoft.opensciencegrid.org cvmfs repository EGI
134042 dteam closed top priority 15/03/2018 29/03/2018 Other Test Ticket. WLCG
134290 cms solved urgent 28/03/2018 28/03/2018 Monitoring T1_UK_RAL failing SAM tests EGI
134309 none verified top priority 28/03/2018 29/03/2018 Other This TEST ALARM has been raised for testing GGUS alarm work flow after a new GGUS release. WLCG
134362 cms solved urgent 01/04/2018 03/04/2018 CMS_SAM tests SAM3 CE and SRM problems at T1_UK_RAL WLCG
134373 ops solved less urgent 02/04/2018 03/04/2018 Operations [Rod Dashboard] Issue detected : org.nordugrid.ARC-CE-SRM-result-ops@arc-ce02.gridpp.rl.ac.uk EGI
Availability Report
Target Availability for each site is 97.0% Red <90% Orange <97%
2018-03-28 100 100 100 100 100 100
2018-03-29 100 100 100 100 100 100
2018-03-30 100 100 100 100 100 100
2018-03-31 100 100 100 100 100 100
2018-04-01 100 100 100 50 100 100 Certificate problem within CMS
2018-04-02 100 100 100 100 100 N/A
2018-04-03 100 100 100 100 100 100
Hammercloud Test Report
Target Availability for each site is 97.0% Red <90% Orange <97%

Key: Atlas HC = Atlas HammerCloud (Queue RAL-LCG2_UCORE, Template 841); CMS HC = CMS HammerCloud

Day Atlas HC CMS HC Comment
2018/03/28 95 99
2018/03/29 96 100
2018/03/30 96 99
2018/03/31 100 100
2018/04/01 100 100
2018/04/02 100 100
2018/04/03 99 100
Notes from Meeting.
  • None yet